# ROSE: A Recognition-Oriented Speech Enhancement Framework in Air Traffic Control Using Multi-Objective Learning

### v1

@@ 4 Can understand ATC instructions clearly and repeat the spoken words
To validate the effectiveness of the multi-objective learning mechanism, ablation experiments are conducted by combining different loss functions in the proposed model (A3). As shown in Table 2 (group B), by incorporating the SC loss into ASR-required features learning, the ASR performance can be considerably enhanced by the multi-objective framework without any model retraining, which validates the core motivation of this work. Compared to the spectrogram features, the MFCC-based SC loss achieves better improvement since it is a more sophisticated handcrafted feature (B1 vs. B2). The best performance is reached by combining the two SC losses, from 4.29% to 3.50% (A10 vs. B3). Except for the ASR performance, the proposed model also harvests higher SE performance. The results can be attributed that the multi-objective loss learns the common representations across the two optimization objectives thanks to its powerful ability on feature learning.
@@
To validate the ability of the proposed attention modules, ablation experiments are conducted by combining different attention blocks in the single-objective model (B3). As can be seen from Table 2 (group C), both ABSF and CSAtt blocks can enhance the model performance in terms of both the SE- and ASR-related metrics. Specifically, compared to the ABSF block, the CSAtt can achieve higher performance improvement by optimizing the learned features, where objective metrics are improved by about 0.23 on average and LER is decreased by 0.05% (C1 vs. C2). To be specific, the attention module focuses on informative feature channels to suppress noise interference, while highlighting selective frames to fit the data distribution between features and text labels. Finally, the proposed model can achieve the best performance by combing the two attention blocks (B3 vs. C3), which validates the proposed model architecture.
@@
Experiments on public datasets
Experiments for the SE
Several representative SOTA models with different inputs are selected to compare with ROSE on Voice Bank + DEMAND dataset (Valentini-Botinhao et al. 2016). Specifically, four types of input are classfied as: 1) Waveform, 2) Magnitude, 3) Real-imaginary (RI) and 4) RI + Magnitude.
The experimental results are reported in Table 3. Although ROSE does not achieve the best performance on COVL and ROSE, it still outperforms comparative SOTA models in terms of other metrics. Inside the MetricGAN (+) methods, due to the PESQ-oriented optimization target involved in the training procedure, the models have achieved cliff-like superiorities on PESQ, while other metrics are inferior. This result shows that the objective-oriented loss function can improve task-oriented feature learning. In addition, it can also be found that the traditional method (Wiener) can achieve better performance on the public dataset with artificial noise models, which also supports the research specificities of the ATC speech corpus with real-world noise and speech echo phenomenon.
Experiments for the ASR
To validate the ASR performance of ROSE on the public dataset, according to the mixed type of signals, two datasets based on test-clean samples in the Librispeech dataset (Panayotov et al. 2015) are artificially synthesized, which are additive noise and simulative echo datasets. Specifically, the mixed noisy samples of the additive noise dataset are formed by mixing each utterance of test-clean samples with the noises from the test set of DNS benchmark (Reddy et al. 2020) with 4 SNR levels (-3, 0, 3, and 6 dB).
@@
Several models with different inputs trained on Voice Bank + DEMAND dataset are selected to enhance blind artificial datasets. Acoustic model based on DS2 retrained from scratch on Librispeech dataset is used to validate the ASR performance.
The experimental results are reported in Table 4. Speech is a continuous sequence, in which there are strong temporal correlations between adjacent frames. Compared to additive noise, signal overlapping severely degrades the discriminability of text-related features, leading to collapsed representatives. Compared with waveform, taking T-F features (Magnitude or RI) as input will inevitably lose some potential linguistic features due to additional operation (e.g. STFT), resulting in inferior ASR performance (The $a$LER of CRN is even worse than that of Mixed). The proposed attention modules and multi-objective mechanism achieve improvements on both two datasets, where the model combined with ASR-oriented target achieves significant results ($a$LER from 19.12% to 12.75% and $e$LER from 29.44% to 22.31%, respectively).
Visualization
@@
Other relevant visualizations are shown in Appendix.
Conclusions
@@
In this study, a recognition-oriented speech enhancement model (ROSE) based on multi-objective learning is proposed to achieve the SE task of the ATC domain, while improving the ASR performance without model retraining. Specifically, The ABSF and CSAtt blocks highlight informative information and suppress redundant factors, enhancing the ability of the model to learn objective-related features. The multi-objective learning acquires the common representations across the two optimization objectives by combining the task-oriented loss function, achieving mutual promotion of SE and ASR tasks performance. The experiment results show that the proposed framework can not only achieve the desired performance on the ATC corpus, but can be generalized to the public open-source datasets.

### v2

@@ V-B Ablation Study
By incorporating the SC loss into ASR-required features learning, the ASR performance can be considerably enhanced by the multi-objective framework without any model retraining, which validates the core motivation of this work. Except for the SE performance, the proposed model also harvests higher ASR performance, which can be attributed that the multi-objective loss learns the common representations across the two optimization objectives thanks to its powerful ability on feature learning.
@@
To validate the ability of the proposed attention modules, ablation experiments are conducted by combining different attention modules in the single-objective model (B6). As can be seen from Table V (group C), both ABSF and CSAtt modules can enhance the model performance in terms of both the SE- and ASR-related metrics. Specifically, compared to the ABSF module, the CSAtt can achieve higher performance improvement by optimizing the learned features, where the first four SE-related metrics are improved by about 0.13 on average, and both DS2-CER and W2V2-CER are decreased by 0.05% (C1 vs. C2). To be specific, the attention module focuses on informative feature channels to suppress noise interference, while highlighting selective frames to fit the data distribution between features and text labels. Finally, the proposed model can achieve the best performance by combining the two attention modules (B6 vs. C3), which validates the proposed model architecture.
TABLE VII: ASR results on two artificial datasets. “$a$” denotes additive noise, “$e$” denotes simulative echo.
Model DS2 W2V2
$a$CER(%$\downarrow$ $e$CER(%$\downarrow$ $a$CER(%$\downarrow$ $e$CER(%$\downarrow$
Mixed 33.36 42.72 7.03 10.22
SE-Conformer 21.43 31.10 6.54 8.12
DCCRN-E 26.69 33.67 6.88 9.72
BSRNN 23.79 32.23 6.73 8.55
ROSE (Ours) 12.75 22.31 4.39 6.55
   $w/o$ $L_{ASR}$ 19.12 29.44 6.29 7.99
   $w/o$ $attn.$ 17.86 26.56 5.57 7.22
test-clean 9.97 2.49
@@
V-C Experiments on public datasets
V-C1 Experiments for the SE
To validate the SE performance of the proposed model on the public dataset, in this section, several representative SOTA models with different inputs are selected to compare with ROSE on Voice Bank + DEMAND dataset [58]. Specifically, five types of input are classfied as: 1) Waveform, 2) Magnitude, 3) Phase + Magnitude, 4) Real-imaginary (RI) and 5) RI + Magnitude.
In the Voice Bank + DEMAND dataset, the clean speeches were recorded from 30 speakers, 28 of which are used for the training set and 2 for the test set. To formulate the noisy speeches in the training set, a total of 10 noise models (2 artificial and 8 from the DEMAND dataset [59]) with 4 signal-to-noise-ratio (SNR) levels (0, 5, 10, and 15 dB) are performed on clean speeches (28 speakers) to generate artificial data pairs to support the SE task. To generate an independent test set, 5 unseen noise models (all from the DEMAND dataset) with 4 SNR levels (2.5, 7.5, 12.5, and 17.5 dB) are performed to clean speeches (2 speakers) to obtain the noisy test set. Finally, the training set and test set are 11572 and 824 speech pairs respectively. All samples are clipped to 4 seconds long and audio samples are made available online [^3] https://github.com/XCYu-0903/ROSE.
For the experimental configurations, the setups of the ROSE are the same as mentioned in Section IV-B. The experimental results are reported in Table VI. Although ROSE does not achieve the best performance on COVL and PESQ, it still achieves optimal performance on CSIG and STOI and sub-optimal performance on CBAK. Inside the MetricGAN-related methods, due to the PESQ-oriented optimization target involved in the training procedure, the models have achieved cliff-like superiorities on PESQ, while other metrics are inferior. This result shows that the objective-oriented loss function can improve task-oriented feature learning. In addition, it can also be found that the traditional methods (Wiener and Spec-Sub) can achieve better performance on the public dataset with artificial noise models, which also supports the research specificities of the ATC speech corpus with real-world noise and speech echo phenomenon.
V-C2 Experiments for the ASR
To validate the ASR performance of ROSE on the public dataset, according to the mixed type of signals, two datasets based on test-clean samples in the Librispeech dataset [60] are artificially synthesized, which are additive noise and simulative echo datasets. Specifically, the mixed noisy samples of the additive noise dataset are formed by mixing each utterance of test-clean samples with the noises from the test set of DNS benchmark [61] with 4 SNR levels (-3, 0, 3, and 6 dB).
@@ V-C2 Experiments for the ASR
Comparative models are trained on Voice Bank + DEMAND dataset to enhance blind artificial datasets (i.e., SE-Conformer, DCCRN-E, BSRNN and ROSE). Two ASR models are applied to evaluate the enhanced artificial set, including a supervised DS2 model and an SSL-based W2V2 model. Specifically, the DS2 model is trained on Librispeech 960-hours training dataset, in which the input feature is the 81-dimension spectrogram. The pre-trained W2V2 model provided on fairseq [^4] The BASE W2V2 model pre-trained on Librispeech 960-hours training set is used. is fine-tuned on Librispeech dev-clean set with the CTC loss function. All the above ASR models are end-to-end models without language models.
The experimental results are reported in Table VII. Speech is a continuous signal with strong temporal correlations between adjacent frames. Compared to additive noise, signal overlapping severely degrades the discriminability of text-related features, leading to collapsed representations [62, 63]. Similar to results in Table III, compared to the T-F domain approaches, the time-domain approaches achieve better performance in additive noise and simulated echo scenarios. The proposed attention modules and multi-objective learning mechanism achieve performance improvements on all datasets, and multi-objective learning can provide higher CER reduction, confirming the generalizability of the proposed technical improvements on different kinds of speech noise.
V-D Visualization
V-D1 Enhancement Results
@@
Refer to caption Refer to caption Refer to caption
(a) Noisy (b) Enhanced (without ABSF) (c) Attention mask of the first layer
Refer to caption Refer to caption Refer to caption
(d) Clean (e) Enhanced (with ABSF) (f) Attention mask of the last layer
Figure 4: Visualization for the ABSF module.
Refer to caption Refer to caption Refer to caption
(a) Channel weights of the first layer (b) Attn. maps of the 38th channel (c) Attn. maps of the 15th channel
Refer to caption Refer to caption Refer to caption
(d) Channel weights of the last layer (e) Attn. maps of the 34th channel (f) Attn. maps of the 39th channel
Figure 5: Visualization for the channel attention of the CSAtt module.
Refer to caption Refer to caption Refer to caption
(a) 1st layer of the Encoder (b) 3rd layer of the Encoder (c) 5th layer of the Encoder
Refer to caption Refer to caption Refer to caption
(d) 5th layer of the Decoder (e) 3rd layer of the Decoder (f) 1st layer of the Decoder
Figure 6: Visualization for the sequence attention of the CSAtt module.
V-D2 The ABSF Module
To validate the performance of the ABSF module, we select an example from the test set and visualize it in the following cases: clean waveform, noisy waveform, enhanced waveform (with ABSF or not), and the attention mask $A_{i}$ (mentioned in Section III-C) of the first layer of the encoder (shallowest) and the last layer of the decoder (deepest). All feature representatives in this work are adaptively adjusted for visualization purposes.
As shown in Figure 4, the final attention mask (f) behaves as learning a voice activity detector that preserves features as much as possible to focus on active regions of the voice. Compared with the result (b) obtained by directly concatenating the hierarchical features in skip connections, the ABSF module (e) improves the feature space of the model, obtaining the enhanced signal that is closer to the clean signal (d). According to (d), (e) and (f), we can find that the learned attention mask is similar to the positive amplitude profile of the clean waveform, achieving that by assigning higher weights to valid voice activity frames and lower weights to redundant frames, further validating the fusion processing between encoders and decoders.
V-D3 Channel Attention of the CSAtt Module
Since spoken words and noise are distributed in different speech feature channels, the purpose of the channel attention of the CSAtt module is to give higher weights to the channels with rich spoken information and lower weights to those with noise.
To confirm the channel attention, we visualize the channel attention weights $W_{C}$ (mentioned in Section III-D) obtained from the channel path of the CSAtt module of the first layer of the encoder and the last layer of the decoder, respectively. We further visualize the attention maps $X_{C}$ corresponding to the channel dimensions in $W_{C}$ that are assigned the highest and lowest weights. The visualization sample is the same as that of in ABSF module (as shown in Figure 4).
As shown in Figure 5, (a) and (d) represent the distribution of channel attention weights in the selected network layers (hidden dimensions is $H=48$), respectively, in which assigned highest and lowest weights of channel indexes are marked in red. (b) and (c) are the attention maps of the assigned lowest weight (indexed as 38) and highest weight (indexed as 15) in (a), respectively. while (e) and (f) are the attention maps of the assigned lowest weight (indexed as 15) and highest weight (indexed as 20) in (d), respectively. In general, in the first encoder layer, (b) and (c) present a similar distribution in the different attention channels, which confuses the model to learn effective spoken information in the shallow layer of the neural network. On the contrary, in the deeper decoder layer, (e) and (f) learn the discriminative attention weights to suppress the noisy features marked by red rectangles, in which the corresponding time-domain samples are also marked in Figure 4 (a) and (d). The visualization also confirms that channel attention achieves the assignment of the expected weight for the SE task.
V-D4 Sequence Attention of the CSAtt Module
The sequence attention of the CSAtt module is required to focus on speech central frames with valid phonemes while filtering marginal echo frames. In the feature domain, the process of sequence attention weighting the frames is the process of continuously fitting the feature maps to the clean distribution.
To intuitively understand the learning process of the sequence attention mechanism, we visualize the sequence attention maps $X_{L}$ (mentioned in Section III-D) obtained from the sequence path of the CSAtt module of the first, third, and fifth encoder layers and their corresponding decoder layers, where the input utterance is the same as that in ABSF visualization.
As shown in Figure 6 (layers from shallow to deep), in the shallow layer (a), we observe that attention maps present a mixture of clean and noisy features, which fail to distinguish informative signals. In (b), the result roughly marks the area where valid information may appear, but it is still not clear. The weights assignment of valid signals is gradually revised from layer (c) to layer (d) until layer (e) achieves convergence, which shows that the ability to reconstruct the clean features is improved with the stacking of sequence attention.
TABLE VIII: Experimental results for AEC and dereverberation tasks.
[…]
Noisy 1.80$\pm$.01 3.66$\pm$.06 2.24$\pm$.02 4.66$\pm$.24
NLMS [64] 1.50$\pm$.03 4.82$\pm$.09 1.55$\pm$.04 12.89$\pm$.46
Align-CRUSE [65] 1.84$\pm$.03 4.07$\pm$.10 1.68$\pm$.03 10.99$\pm$.38
NKF [66] 1.93$\pm$.03 3.89$\pm$.09 1.93$\pm$.04 10.52$\pm$.36
WPE [67] 1.55$\pm$.04 4.77$\pm$.10 1.61$\pm$.04 11.67$\pm$.39
WD-TCN [62] 2.78$\pm$.04 2.65$\pm$.03 4.04$\pm$.04 6.09$\pm$.27
RVAE-EM-S [63] 2.81$\pm$.04 2.61$\pm$.02 4.23$\pm$.03 5.11$\pm$.25
ROSE (Ours) 3.09$\pm$.03 2.12$\pm$.05 5.33$\pm$.19 3.50$\pm$.18
TABLE IX: Experimental results of integrating the proposed ABSF, CSAtt and $L_{ASR}$ into SE-Conformer.
[…]
3 3.11$\pm$.04 4.05$\pm$.19
V-E Discussions
V-E1 Confirmation of the specific natures of the speech echo in the ATC domain
As mentioned in Section I, the type of signal that causes the speech echo cannot be defined with certainty. Therefore, we additionally use acoustic echo cancellation (AEC) methods and dereverberation methods to explore the effectiveness of different signal processing methods in eliminating the speech echo. The clean-noisy pair dataset is used to conduct relevant experiments.
For the AEC methods, three open-source baselines are selected to conduct the experiments, including a traditional method NLMS [64] and two DNN-based methods Align-CRUSE [65] and NKF [66]. According to previous work [6], we can only collect the received speech (clean) and the combined (noisy) speech in a real-world ATC environment. To simulate the far-end signal and not introduce other signals, we generate the simulated sent speech only by applying random time offsets between 0 and 200 milliseconds to the received speech and then clip it to ensure the temporal alignment with the original signal. Therefore, to fit the AEC scenarios, we regard the received speech as the near-end signal, the simulated sent speech as the far-end reference signal, and the combined speech (with the speech echo) as the mixed signal. The purpose of the AEC task is to remove the far-end signal from the mixed signal to obtain the estimated near-end signal, which is also the same as the purpose of the SE task. For dereverberation methods, three open-source baselines are selected to conduct the experiments, including a traditional method WPE [67] and two DNN-based methods WD-TCN [62] and RVAE-EM-S [63]. To fit the dereverberation scenarios, we regard the received speech as the clean signal and combined speech as the reverberant signal.
For the metrics of the above tasks, we adopt the objective evaluation metrics on the test set, including PESQ [51], cepstral distance (CD), speech-to-reverberation modulation energy ratio (SRMR) [68] and DS2-CER to measure effects of the signals. Except for SRMR, the calculation of other metrics requires the corresponding reference speech. For CD and DS2-CER, lower values indicate better performance; for other metrics, higher values indicate better performance. We also report the mean results and their 95% confidence interval.
As shown in Table VIII, the AEC-related methods not only fail to remove the speech echo, but also lead to worse speech quality. The reason is that in the common AEC scenarios, the far-end and near-end signals are different signals, and AEC methods are able to achieve signal separation based on the reference signals. In the ATC scenarios, the far-end and near-end signals are highly similar (almost the same, and all signals are issued by the same ATCO), and the AEC methods are unable to capture effective features to support SE and ASR tasks. Similar to the traditional methods of SE, the WPE model fails to remove the speech echo and even obtains worse performance than noisy speech. Although the DNN-based dereverberation methods can obtain the desired SE performance due to the task objective with the SE, they still suffer from poor ASR performance even compared to the noisy speech, which fails to meet the requirements for both the SE and ASR tasks in this work. Therefore, the experimental results confirm that the speech echo is a specific mechanism that cannot be removed by common AEC and dereverberation methods, which also supports the motivation of our proposed ROSE framework.
V-E2 Exploration for the Generalizability of the Proposed Technical Improvements
Previous experimental results show that all the proposed technical improvements (i.e., ABSF module, CSAtt module and ASR-oriented loss function) achieve the desired performance on the architecture used in this work. As a similar U-Net architecture, the SE-Conformer model is also considered to explore the generalizability of the proposed modules. Specifically, the ABSF module is integrated into skip-connections and the CSAtt module is integrated to the end of each encoder layer and the beginning of each decoder layer, respectively. The ASR-oriented loss function is implemented by the multi-objective learning mechanism.
As shown in Table IX, the number in the CSAtt column indicates the order of layers in which the CSAtt module is integrated into the encoder/decoder. For example, option 1 indicates that it is integrated only at the first layer of the encoder and the last layer of the decoder, and option 5 indicates that it is integrated at every layer of the encoder/decoder. The experimental results show that when ABSF, CSAtt, and $L_{ASR}$ are integrated into the SE-conformer, it achieves desired improvements on both SE and ASR tasks. In addition, the specific configurations of the CSAtt module are also determined by designed experiments. The results confirm the effectiveness of multi-objective learning for the task in this work with different model architectures. Most importantly, the results also prove the performance improvement of the ABSF, CSAtt and $L_{ASR}$ on both the CNN- and Conformer-based U-Net architecture, which confirm the contributions of all the proposed technical improvements in this work.
VI Conclusions
@@
In this study, a recognition-oriented speech enhancement model (ROSE) based on multi-objective learning is proposed to achieve the SE task in the ATC domain, which enables us to improve the ASR performance without model retraining. Specifically, the ABSF and CSAtt modules highlight informative information and suppress redundant factors, enhancing the ability of the model to learn objective-related features. The ASR-oriented loss function is designed to guide the model to capture ASR-related representations in spectral scales. The multi-objective learning acquires the common representations across the two optimization objectives by combining the task-oriented loss function, achieving mutual promotion of SE and ASR tasks performance. The experiment results show that for both the SE and ASR tasks, the proposed framework can not only achieve the desired performance on the ATC corpus, but also be generalized to the public open-source datasets. The proposed framework can generate high-quality speech to address the speech echo in the ATC domain, as well as preferred ASR performance without any model retraining. In addition, all the proposed technical modules contribute to the expected performance improvements, as demonstrated by ablation studies. These improvements can also be applied to other neural architectures to enhance task performance. The visualization results showed that the proposed attention mechanism can learn informative weight distributions, providing interpretability for its performance effectiveness.